A TOUR ON THE LIFE AND WORK OF A. M. NAVEIRA (Part II)

Angel Ferrández (Universidad de Murcia)

Profesor Naveira, Antonio; Tatá; Marta; Bruno; Marcos;

Queridos amigos/as, mesdames et messieurs, ladies and gentlemen.

Soy muy feliz de estar aquí y tener el privilegio de impartir esta charla. La tarea me debería resultar fácil, por tratarse del Prof. Naveira, mi maestro y amigo Antonio, y, sin embargo, me preocupa saber transmitir una serie de hechos absolutamente necesarios para entender el por qué del fenómeno Naveira y el por qué de aprecios incondicionales como el mío. Sólo Vdes., al final, me dirán si he acertado.

I’m very happy to give this lecture. It should be quite easy for me, because Professor Naveira is my teacher and friend. However, I’m not sure to be able to transmit to you, to get into your souls, a series of absolutely necessary facts to understand why the Naveira phenomenon and why many of us love him. At the end of this talk, please, let me know if I have got my objective.

Estoy seguro que cualquiera del Comité Organizador habría sido perfectamente capaz de disertar sobre la obra del Prof. Naveira, pero al final hemos sido elegidos los que menos hemos trabajado para que este evento llegase a buen puerto. En nuestra descarga, daremos todo lo mejor de nosotros. También es cierto que quizás hemos intimado más con él y su familia. Por eso, y porque somos de los discípulos más viejos, que no más aventajados, se nos permitirán ciertas licencias.

I’m sure that anyone of the Organizing Committee was able to prepare a very nice talk on this subject. However, Prof. Hervella and myself have been chosen according to the following three reasons: (i) we’re who less time have spent to organize this event; (ii) we’re probably the closest to Antonio and his family; and (iii) we’re Naveira’s oldest, but not the best, disciples. Be sure that we’ll do our best to reach our goal.

Cuando el Prof. Hervella y un servidor propusimos distribuirnos la tarea de exponer la obra de Antonio, yo le cedí cortésmente el turno a Luis, quien con su maestría innata, casi sin darle importancia, me espetó: “Yo hablaré de la obra de Naveira hasta las casi-producto”. Acepté de inmediato, pero no tardé más de dos minutos en darme cuenta de que acababa de robarme la cartera. La finísima inteligencia del Prof. Hervella se había puesto una vez más de manifiesto, cosa sólo posible en la mente de un matemático dedicado a la Geometría Diferencial.

When Prof. Hervella and myself started with the organization of this talk, I gently yield the turn to him, who quickly said to me: “I’ll lecture on Naveira’s work until almosthermitian and almost-product manifolds”. In a few minutes I felt that my wallet had been stolen. Once more, Prof. Hervella showed to me he has a very quick mind which is characteristic of people working on Differential Geometry and related topics.

En todo caso, mi decisión estaba tomada en el sentido de buscar principalmente la amenidad, que no significa renunciar a la seriedad, para explicar el excelente trabajo de una persona que se ha entregado en cuerpo y alma a una tarea en la que cree profundamente, pero con ese grado de humanidad que lo hace único, singular e irrepetible. Antonio, afortunadamente, rompe moldes. Resaltaré una serie de hechos clave en su vida que lo caracterizan y lo hacen entrañable.

I decided to look for amenity, keeping of course seriousness, to explain the excellent work of Prof. Naveira, which is nothing but the answer of someone devoted himself to studying; someone always having blind faith of his doing. And everything done with a bit of geniality that makes him unique, singular and unreproducible. Antonio, as Spain, is different. I’ll point out a series of key facts all through his life which characterize Prof. Naveira and become him a great guy.

Decir que es mi maestro, siempre mi amigo, es algo que para mi representa tanto que lo proclamaré mientras viva y me quede una pizca de voz. Y a él le dediqué, le dedicamos, cuando cumplió sus 60 años, el mejor trabajo de nuestra vida, como no podía ser de otra manera.

I always like to name him El Maestro (The Teacher), and this fact is so important for me that I’ll forever proclaim it provided that I keep a little bit of breath. On ocassion of his 60 birthday, I, we, dedicated him our best paper.

Véase TRANS. 1: Artículo de Barros, Ferrández y Lucas en Nuclear Physics

See TRANS. 1: Paper of Barros, Ferrández and Lucas in Nuclear Physics

La labor científica del Prof. Naveira está ahí, en cualquier momento puede ser consultada. La importancia de la misma habla por sí sola, de tal manera que no estaríamos aquí, celebrando este congreso, si no fuera por la categoría del personaje y su obra. Pero coincidirán conmigo en que, paralelamente, Antonio llevó a cabo una labor humana, de creador de escuela, tanto o más importante que aquella, y que para mi constituye la esencia de este simposio.

The scientific work of Prof. Naveira is well known and anyone have enough tools to consult it. It is so important that I’m sure that we want to spend some of our time because we believe it is our duty to do it. Furthermore, I think you will agree with me, the main project carried out by Antonio was the creation of a Differential Geometry School. This is, probably, the very reason of this meeting.

Si bien es cierto que este homenaje es un pretexto para hablar de Geometría Diferencial, también lo es que, si sólo fuese para eso, probablemente quien les habla no estaría aquí. He tenido mucho que ver con esta celebración, al menos con la inspiración y la filosofía de la misma, cosa que deseo dejar patente, pues dos de los principios que mueven mi vida radican en alegrarme de los verdaderos éxitos de mis prójimos y en aplicarme el dicho aquél de Sancho “es de bien nacido ser agradecido”. Así pues, más que ponderar la obra de Naveira del último decenio, destacaré una serie de detalles, momentos, frases y anécdotas que nos acercarán al perfil humano del Prof. Naveira, lo harán más comprensible y quizás así puedan entender el por qué de ciertas adhesiones incondicionales. Es muy posible que no llegue a alguno de Vdes., pero quienes me entiendan llegarán a apreciar de verdad la personalidad de Antonio. Esta es su semana y me gustaría que nada pudiese turbar su felicidad a lo largo de ella.

The fact is that this tribute to Prof. Naveira is a pretext to talk about some recent advances in Differential Geometry. However, if this were the only aim of this meeting, I confess that probably I would not have attended it. I strongly bet for this celebration, its inspiration and philosophy, because two, among the principles governing my life, are, on one hand, to be happy with the success of my friends; and, on the other hand, to apply myself that saying of Sancho to Don Quijote (Don Quichote): “It is only well born who is able to give thanks” (free translation). So then, besides to speak highly of Naveira’s work in the past ten years, I’ll point out some details, moments, sentences and anecdotes that we’ll bring ourselves near to the human profile of Prof. Naveira, so that you can find him more understandable and you can then understand why certain unconditional adhesions. The core of my speech is going to anybody that want to know the very personality of Antonio. We wish him, and his family, a very happy week.

Conocí a Antonio en septiembre de 1975, en la Universidad de Granada. Unos cuantos jóvenes, de la misma promoción, cargados de ilusiones, acabábamos de incorporarnos al Departamento de Geometría y Topología con la única idea de hacer una tesis en Geometría, cosa poco menos que imposible por entonces. En nuestro primer contacto con Naveira quedamos asustados, lo vimos como un extraterrestre cuando nos dimos cuenta que llegaba el primero por las mañanas, se iba el último por la noche, se tomaba un café en tres minutos, comía en apenas una hora. En fin, pensamos, tenemos muchas ganas de investigar, pero con personas normales. Claro está, nosotros llegábamos a una hora “prudente”, las 11 a.m., como no podía ser de otra manera, pues la noche anterior era de obligado cumplimiento cerrar todos los chiringuitos del Campo del Príncipe. El pobre de Antonio estuvo a punto del suicidio, pues solicitaba, amargamente, algún joven a quien dirigir la tesis, ya que llegaba con un montón de problemas recientes y, tal era su generosidad, casi resueltos. Pero, créanme, daba miedo, tanto por el ritmo de trabajo que adivinábamos, como por aquellas lindezas de Kaehlerianas, nearly, almost, quasi, gallegas, don Enrique, Gray, Vanhecke, Santiago, París, Lovaina, Maryland,...

I met Naveira in September 1975 in the University of Granada. A few young men of the Department of Geometry and Topology had decided to study Geometry in order to make a thesis. It seemed, at that time, an impossible task because we had no direction, no open problems, no objectives, no ..., no ...; however, we found a never open pair of books: K-N I & II. The first time we talked about working with Prof. Naveira we spoke different languages. Otherwise, we thought he were coming from another world. He arrived, to the department, the first, at 8 a.m.; he leaved it the latter; he had a cup of coffee in 3 minutes; he took lunch in 45 minutes; and so on. As a matter of fact, we cannot arrive until 11 a.m., because we had to close all pubs in the Prince Field. Antonio was seeking for any young mathematicien desesperately to start a thesis. He came to Granada with a lot of open problems, and almost solved. But, please, believe me, we were afraid of the way working of Antonio and such an amount of strange things such as Kaehlerian, nearly, almost, quasi, G_{1} and G_{2}, don Enrique, Gray, Vanhecke, Santiago, París, Lovaina, Maryland,...It seemed we were in Hollywood.

Aunque todos ansiábamos empezar a trabajar en nuestras tesis, quiero destacar un hecho crucial para mi. Se trataba de respetar un pacto de caballeros, a todas luces inquebrantable: Manolo Barros tenía reservada la primera plaza. Así que Antonio, que tardó en entenderlo, tuvo que esperar, pero pronto se percató de que, aunque algo gandules en las formas, teníamos un fondo de seriedad que rápidamente captó y enseguida hizo suyo. En enero de 1976 el Prof. Barros regresó de sus deberes militares y en apenas un año, aquellos problemas que traía casi resueltos quedaron extraordinariamente bien resueltos y publicados, dando lugar a la primera tesis de Naveira allende Galiza. Fue la tesis del Prof. Barros, finalizada en la Univ. de Valencia, pero leída en la de Granada a mitad de diciembre de 1976. Este hecho merece una mención especial, pues a raíz de esta fecha se inicia, bajo la dirección del Prof. Barros, la historia reciente y brillante del grupo de Geometría de la Univ. de Granada. Y tras las huellas de Manolo, allá que nos fuimos A. Ramírez, A. Ferrández y F. J. Carreras. El paso del tiempo ha confirmado el acierto de nuestra decisión.

As a matter of fact, we yearned for working in our thesis. However, I wish to point out a key fact: we had an unbreakable agreement, accordingly Manolo Barros should be the first one to begin a thesis with Antonio. It was a little difficult to explain it to Antonio, since he had not yet met Barros. In January 76, Manolo came back from military duties and hardly a year those open problems, that Antonio brought himself, were nicely well solved. This was the first thesis of Naveira out of Galiza, which finished in this university, althought its reading took place in Granada about the middle of December 76. It deserves a special mention, since this is the root, under the direction of Prof. Barros, of the recent and brilliant history of the Granada Differential Geometry team. Following Barros’ tracks, A. Ramírez, F. J. Carreras and myself went to Valencia looking for a promising future. The time has been said to us that we were right.

No es este el momento de alabar las excelencias de la investigación en Geometría en la Universidad de Granada, pero no me queda más remedio que decir que la llegada de Naveira a Granada fue un hecho histórico, pues a partir de entonces se debe hablar de “Granada antes de Naveira” y de “Granada después de Naveira”.

I do not pretend to talk about Granada University, but Naveira is directly concerned with the success of the current Differential Geometry group. I like to speak of “Granada before Naveira” and “Granada after Naveira”.

Pero ¿cómo aterrizó Naveira en Granada? ¿En un Boeing? ¿En un Airbus? Que va; ¡se admiten apuestas!

How did Naveira land in Granada? In a Boeing 747? In an Airbus? Lay a bet!

Véase TRANS. 2: El Purísimo  See TRANS. 2: The Purísimo

Antonio cambió radicalmente nuestras costumbres. Pero también supo aprender de nosotros: yo asistía a sus clases, tomaba apuntes, los ponía en limpio y luego él los hacía suyos. Él nunca tuvo apuntes personales, excepto los que tomó de las magistrales clases de Deheuvels en París; que por cierto tiramos a la basura un desgraciado día. Para curvas y superficies seguía, como no, el libro de D. Enrique Vidal.

Antonio changed radically our way of life. However, he also learnt from us. I attended his classes, took personal notes, put in a clean and easy to read form and he asked me for them. It was for me a great pleasure. He never used to have personal notes to teach, up to those taken from Prof. Deheuvels classes in Paris. They have a sad history: we need them for my thesis and a bad day, cleaning people throw them out the rubbish.

Otra de sus genialidades fue la de contratar unas clases de inglés, pero con un nativo tartamudo. Al cabo de dos semanas ya éramos capaces de decir “Yo soy un hombre = I am an egg”. Aparte de lo “acertado” de la traducción, pensábamos que la doble “g” era producto de la tartamudez de nuestro “profe”.

We found Antonio’s genialities anywhere and anytime. For instance, he took a very special English teacher: a stuttering teacher. After two weeks, we had to answer a difficult quiz: “How do you say in English Yo soy un hombre”. One of us gave the answer: I am an egg. The very joke is not really de answer, but the double “g” of “egg”, because we were sure that the second “g” was produced by the stuttering of the teacher. Incredible!

Su curso 75-76 de Curvas y Superficies se hizo archifamoso cuando introdujo las “superficies compatas”. Desde entonces, es muy conocida la “Clasificación de Naveira de las superficies compatas”, que queda como sigue:

The academic year 75-76 was included in the annals of the University of Granada. Naveira’s classroom was crowed, because he succeeded introducing “the compat surfaces” (“compat” should be translated by “with legs”; maybe, a little difficult for non Spanish speaking people). Anyway, “Naveira’s classification of compat surfaces” is as follows

Véase TRANS. 3: Los modelos de “superficies con patas”

See TRANS. 3: The models of “surfaces with legs”

Su frase favorita por entonces era “Yo introduje la curvatura en España”. Claro, uno está tentado a imaginarse

A famous sentence: “I introduced the curvature in Spain”. Indeed, it is easy to imagine the following

Véanse TRANS. 4 (España antes de Naveira) y 5 (España después de Naveira)

See TRANS. 4 (Spain before Naveira) & 5 (A piece of Spain after Naveira)

A través de un curso de doctorado, Antonio nos metió de bruces en las variedades casi-hermíticas. Él se manejaba como pez en el agua en este campo y pronto nos contagió. Pero, claro, la nota simpática no tardó en aparecer, pues de tanto “subir y bajar la J”, en nuestro argot hablábamos de “bailar la J”. Sin duda, los mejores coreógrafos del grupo de baile han sido Gray y Hervella.

We followed a postgraduate course where Prof. Naveira taught us the first classification of almosthermitian manifolds. There he felt completely at home and very soon we were able to get up and down with the almost-complex structure, i.e., with the familiar J. In such a way that in our argot we used to say “to dance the J”, like a quite typical Spanish dancing. It is well known that the best choreography of the J-dancing has been, without doubt, carried out by the couple Gray-Hervella.

Me gustaría que mi charla refleje mi aprecio por la persona que me puso en el camino de la luz y como mi mejor homenaje a su trabajo. Pero él mismo sabe que el mejor reconocimiento a la labor de toda una vida no son sólo palabras, sino el haber entendido su filosofía y haberla sabido aplicar creando nuevas escuelas y mostrándonos orgullosos de que cada discípulo haya podido superar al maestro. Y a los hechos me remito, pues sé de la inmensa satisfacción del Prof. Naveira al contemplar el altísimo nivel logrado por la Geometría Diferencial en España y, concretamente, por las escuelas que nacieron a raíz de aquél afortunado encuentro en la Universidad de Granada en septiembre de 1975. Discípulos directos son M. Barros, A. Ramírez, V. Miquel, A. Ferrández, O. Gil, F. Carreras, A. H. Rocamora, S. Segura y X. Gual. “Sus hijos”, en palabras de Antonio. Aquella semilla eclosionó con tal fuerza que hoy puede estar más que orgulloso de sus frutos con “nietos” y “biznietos” en las universidades de Santiago, Coruña, País Vasco, Castellón, Valencia, Politécnica de Valencia, Alicante, Murcia, Politécnica de Cartagena y Castilla-La Mancha. Como prueba de ello, basta decir que, sólo de los “hijos”, se pueden contar, en 25 años, más de 200 artículos en las revistas más prestigiosas.

I’ll be very happy if, through my talk, I have been able to show you that Prof. Naveira is a person that I hold in the highest regard. This is mi little tribute to a man of whom I learned the way to do research and the very spirit of the university. He knows very well these are not only words, but also the understanding of a philosophy and its application. That means that the old school has grown year out and we are very proud to see how the young researchers outshine the teachers. It is well known the selfsatisfaction of Antonio talking about his “sons”, because the Spanish Differential Geometry is going through a sweet period and all of them have to do with. This is a short list: M. Barros, A. Ramírez, V. Miquel, O. Gil, F. Carreras, A. H. Rocamora, S. Segura, X. Gual and myself. (L. M. Hervella, of course, completes the list). The fruits of Granada’s seed yield to second and third generations, so that Naveira has “grandsons” and “great-grandsons” in the universities of Santiago, Coruña, País Vasco, Castellón, Valencia, Politécnica de Valencia, Alicante, Murcia, Politécnica de Cartagena, Castilla-La Mancha and, of course, Granada. I’m very proud to tell you that, in the las 25 years, “Naveira’s sons” have been written more than 200 papers, published in the first ranked journals.

Véase TRANS. 6: (La herencia de Naveira)

See TRANS. 6: (Naveira’s heritage)

¿Por qué se dedicó al estudio de la Geometría Integral? Nunca lo comenté con Naveira, pues coincidió con mi marcha a la Universidad de Murcia. Pero, conociéndole, creo que le movieron dos razones de peso: una, Santaló y su gran personalidad, y su libro; otra, a medida que nos cargamos de años, parece que sentimos la necesidad interior de buscar aplicaciones a esa investigación básica que durante tanto tiempo nos ha ocupado.

Why Naveira engaged in studying integral geometry? I never talked about this question with Antonio, but I know him and he was probably guided by, on one hand, the great personality of Santaló, and his book; on the other hand, according time flies, one needs to do something special, such as to see an application of your early own research.

Un concepto definido sobre los cuerpos convexos, como es el de medidas seccionales promediadas (MSP), “Quermassintegrale”, introducido por Minkowski, será la base a partir de la cual se definen las curvaturas totales, concepto muy importante en el que se fundamenta parte de la reciente investigación desarrollada por Naveira en el campo de la Geometría Integral. Si se imponen ciertas condiciones de diferenciabilidad a la frontera ∂K del conjunto K, se pueden expresar las MSP de K como función de la integral sobre ∂K de las funciones simétricas elementales de las curvaturas principales. Al dar una expresión local de las curvaturas totales aparece una relación entre éstas y la fórmula del volumen de un tubo alrededor de una subvariedad. Este hecho motivó en Naveira el estudio de fórmulas para el volumen de tubos alrededor de ciertas subvariedades en espacios simétricos compactos. Naveira y sus discípulos han aportado interesantes contribuciones en estos campos y pueden encontrar sus artículos en revistas tales como Revue Roumaine, Rendiconti di Matematica, Annals of Global Analysis and Geometry, Diff. Geometry and its Applications, Archiv der Mathematik, Mathematica Scandinavica y Publ. Math. Univ. Debrecen, entre otras.

The notion of averaged sectional measures (ASM), defined on convex bodies K, also known as “quermassintegrales”, given by Minkowski, yields to total curvatures. The recent work of Naveira, in the field of integral geometry, is based on that notion. By assuming certain differentiablility conditions on ∂K, we find the ASM of K in terms of the integral over ∂K of the elementary symmetric functions of principal curvatures. By giving a local expresion of the total curvatures, we find a relationship among them and the formula of the volume of a tube around a submanifold. This is the key fact that yielded to Naveira to studying formulas for the volume of tubes around certain submanifolds in compact symmetric spaces. Naveira and his students have made interesting contributions in these fields which you can find in papers published in Revue Roumaine, Rendiconti di Matematica, Annals of Global Analysis and Geometry, Diff. Geometry and its Applications, Archiv der Mathematik, Mathematica Scandinavica and Publ. Math. Univ. Debrecen, among others.

Están muy de moda las aplicación de la Geometría Integral a la biología y la mineralogía. La rama en la cual se aplican las técnicas de Geometría Integral recibió, en 1961, el nombre de Estereología, entendida como “un conjunto de métodos para la exploración del espacio tridimensional a partir del conocimiento de secciones bidimensionales o proyecciones sobre planos”. Es decir, se trata de una extrapolación del plano al espacio. Uno de los más destacados investigadores en el campo de la Estereología es L. M. Cruz Orive (quien ha estado más de 20 años como investigador en la Universidad de Berna y actualmente es catedrático en la Universidad de Cantabria), con quien Naveira comparte proyectos de investigación desde hace varios años, en concreto desde el año 1991, aunque los dos coincidieron en su etapa de servicio militar sin ellos saber que años más tarde compartirían inquietudes geométricas.

It is in fashion the application of integral geometry to biology and mineralogy. The branch where integral geometry applies has been called Stereology, defined as “the set of methods to explore the 3-space by means of the knowledge of 2dimensional sections or projections over planes”. There is nothing but an extrapolation from plane to space. The number one in stereology is Prof. L. M. Cruz Orive (University of Cantabria), an old friend of Prof. Naveira, who spent more than 20 years as a researcher in the University of Berna. They shared, a few years ago, military duties and now, since 91, they also share applied geometry projects.

Véanse las fotos con el Prof. Santaló

See pictures with Prof. Santaló

Su último éxito: la resurrección de la Real Sociedad Matemática Española.

His last, but not least, success has been the resurrection of the Spanish Mathematical Society.

RESUMEN. Naveira es intuición, ilusión, inquietud, fe, genialidad, empuje. Es infatigable y desconoce el desaliento. Es fiel, es maestro, es amigo. Tiene devoción por sus maestros y antiguos amigos: D. Enrique, Profesores Deheuvels, Lichnerowicz, Gray, Vanhecke, Bourguignon, Aubin. Y le encanta todo “lo francés”. Profesor Naveira, Antonio, que cumplas muchos años y que todos los que hemos tenido el placer de acompañarte los disfrutemos contigo. Muchas gracias por su amable atención.

SUMMARIZING. Naveira is intuition, hopefulness, faith, geniality, energy. He is tireless and does not know the discouragement. He is faithful, a teacher and a friend. He is very fond of his teachers and old friends: D. Enrique, Professors Deheuvels, Lichnerowicz, Gray, Vanhecke, Bourguignon, Aubin. He is also very fond of France and French people. Antonio, have a long life and let us enjoy it with you. Thank you for your kind attention.